My daughter is just learning fractions in 4th grade and I am trying to help her. Unfortunately it has been some time since I have done this. I can't seem to remember how to find the value of x in a fraction...like this x/8 = 1/4.
can you help me?

Answer by stanbon(75887)

x/8 = 1/4
-----
We have to be careful when it comes to
solving problems for young students.
-----
One way to work the problem is to
convert 1/4 = 2/8
---
Then x/8 = 2/8
---
So x = 2
-----------------
Another way to work the problem is to
multiply both sides of the equation
by 8:
8(x/8) = 8(1/4)
---
After canceling you would get:
x = 8/4
--
Then x = 2
==================
Let your daughter decide which method looks
like the way she is being taught.
===================
Cheers,
Stan H.